A rectangular garden is to be surrounded by a walkway of constant width. The garden's dimensions are 30 ft by 40 ft. The total area , garden plus walkway, is to be 1800 ft^2. What must be the width of the walkway to the nearest thousandth?
Thank you.

Answer by Paul(988)

Let the width be x
Since 2 width on each side we represent this as:
30-2x and 40-2x
(30-2x)(40-2x)=1800
1200+2x^2+10x=1800
2x^2+10x-600=0
x^2+5x-300=0
Factor:
(x+20) and (x-15)
x=15
Hence, the uniform width around the garden is 15cm each.
Paul.